Lorie Metals Corporation makes two products, titanium Rims and Posts. Data regarding the two products follow:

   

Direct
Labor-Hours
per unit
Annual
Production
Rims 0.80 29,000 units
Posts 0.80 73,000 units

Additional information about the company follows:

  1. Rims require $20 in direct materials per unit, and Posts require $18.
  2. The direct labor wage rate is $16 per hour.
  3. Rims are more complex to manufacture than Posts and they require special equipment.
  4. The ABC system has the following activity cost pools:
Estimated Activity
Activity Cost Pool Activity Measure Estimated
Overhead
Cost
Rims Posts Total
Machine setups Number of setups $ 34,310 100 100 200
Special processing Machine-hours $ 113,150 1,000 0 1,000
General factory Direct labor-hours $ 1,218,000 4,000 83,000 87,000

Compute the activity rate for each activity cost pool. (Round your answers to 2 decimal places.)

Activity Rate per setup Activity Cost Pool Machine setups Special processing General factory per MH per DLH
Answer #1

Activity rate = Total verhead / Number of activity

Machine setup = 34310 / 200 = 171.55

Special processing = 113150 / 1000 = 113.15

General factory = 1218000 / 87000 = 14.00